OpenStudy (anonymous):

Find the perpendicular slope 3x + 2y=-7 (1,1) Use slope intercept form please!!!

OpenStudy (anonymous):

Okay, so you have to get the y alone. Subtract "3x" from each side: 3x + 2y = -7 -3x -3x And it will look like: 2y = -3x - 7 You have to divide by 2 because remember: "Get the y alone" 2y divided by 2 is of course 1. -3 divided by 2 would have to turn into -3/2x -7 divided by 2 turns into -7/2 *Perpendicular slopes have a negative reciprocal slope. So -3/2 turns into 2/3

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

No problem :D